OpenStudy (xanthe):

What would you expect to observe if hydrogen selenide were passed into PBNO3 solution?

OpenStudy (abb0t):

electron transfer?

OpenStudy (xanthe):

really?I hv no idea yo><

thomaster (thomaster):

i'd expect a black lead(II)sulfide precipitate

OpenStudy (xanthe):

is it NO2 gas?

thomaster (thomaster):

\(H_2S + Pb(NO_3)_2 = PbS(s) + 2HNO_3\)

OpenStudy (xanthe):

Opps sorry u cant observe a gas-.-

thomaster (thomaster):

you can observe a gas :P its bubbles

OpenStudy (xanthe):

Ohyah my bad><

OpenStudy (xanthe):

@thomaster y didnt i think of that?

OpenStudy (xanthe):

>< My brain..

OpenStudy (xanthe):

I cant thank you enough!:D

thomaster (thomaster):

I see i made a mistake i think... \(H_2S + PbNO_3 = PbS(s) + H_2NO_3\)

thomaster (thomaster):

u sure its not Pb(NO3)2?

OpenStudy (xanthe):

But isnt it HSe or H2Se?

OpenStudy (xanthe):

Yes that one!Sorry mines was wrong><

OpenStudy (xanthe):

Its Pb(NO3)2 my bad

thomaster (thomaster):

no selenium has 6 electrons and it needs 8 to compelte its outer shell so it needs to share 2 electrons. Hydrogen has only 1 electron to share so 2 hydrogen atoms share 1 electron with selenium